# Multiple Imputation for Nonresponse in Complex Surveys Using Design Weights and Auxiliary Margins

### v4

@@ 3.2 Variables with Margins for Unit Nonrespondents
We impute each $x_{i2}^{*}$ given $x_{i1}^{*}$ for units with $U_{i}=1$ using a random draw from the adjusted probability mass function $\{\tilde{p}_{i2c}:c=1,\dots,m_{2}\}$, where $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}$ for $c=1,\dots,m_{2}-1$ and $\tilde{p}_{i2m_{2}}=1-\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}$. If $\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}>1$, we compute (3.3) for $c=1,\dots,m_{2}$ and set $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}/\sum_{c=1}^{m_{2}}f_{2c}{p}_{i2c}$. This process ensures the completed-data Horvitz and Thompson (1952) estimates approximately match the sampled $(\hat{T}_{21},\dots,\hat{T}_{2m_{2}})$ in expectation, while also incorporating relationships between $X_{2}$ and $(X_{1},W)$ implied in the working probabilities. We independently repeat the estimation and imputation process in each completed dataset.
When $\{p_{i2c}\}$ derives from a logistic regression of $X_{2}$ on $X_{1}$ (and possibly $W$), using $\{\tilde{p}_{i2c}\}$ is equivalent to adjusting the intercept in that estimated regression, leaving other coefficients alone. Essentially, this implies a model where the log-odds for $X_{2}$ are the same for unit respondents and nonrespondents, per the hybrid missingness MD-AM assumption in (2.2) of no interaction between $U$ and $X_{2}$.
When disregarding $W$ in the logistic regression for $X_{2}$ on $X_{1}$—so that within any completed dataset, $p_{i2c}=\Pr(X_{2}=c|x_{1}=d,U=1)$ for all unit nonrespondents with $x_{i1}^{*}=d$ in that dataset—we can find $\{\tilde{p}_{2c}\}$ by solving a system of equations implied by (2.2). Specifically, we encode the constant log-odds assumption as a set of $(m_{1}-1)(m_{2}-1)$ equations. For $c=2,\dots,m_{2}$ and $d=2,\dots,m_{1}$, we have
$\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=1)}{\Pr(X_{2}=1|X_{1}=d,U=1)}\right)-\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=1,U=1)}{\Pr(X_{2}=1|X_{1}=1,U=1)}\right)$
$=\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=d,U=0)}{\Pr(X_{2}=1|X_{1}=d,U=0)}\right)-\text{log}\left(\frac{\Pr(X_{2}=c|X_{1}=1,U=0)}{\Pr(X_{2}=1|X_{1}=1,U=0)}\right).$ (3.4)
We can estimate the conditional probabilities for units with $U_{i}=0$ via survey-weighted ratio estimators using the completed data; for example,
@@
We also encode the condition (2.5) on $(\hat{T}_{21}^{*},\dots,\hat{T}_{2m_{2}}^{*})$ as a set of $m_{2}$ equations. We have
$\sum_{d=1}^{m_{1}}\sum_{i\in\mathcal{S}}\Pr(X_{2}=c|X_{1}=d,U=1)w_{i}I(x_{i1}^{*}=d,x_{i2}^{*}=c,U_{i}=1)=\hat{T}_{2c}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i2}^{*}=c,U_{i}=0).$ (3.6)
In total, we have $m_{1}(m_{2}-1)$ equations to solve for the $m_{1}(m_{2}-1)$ conditional probabilities, $\{\Pr(X_{2}=c|X_{1}=d,U=1):c=2,\dots,m_{2};d=1,\dots m_{1}\}$, which we use for imputation.
To illustrate, we present the computations for (3.4) and (3.6) for binary $X_{2}$ and $X_{1}$. We use these in the simulations of Section 4. First, we compute $\Pr(X_{2}=1|X_{1}=1,U=0)$ and $\Pr(X_{2}=1|X_{1}=0,U=0)$ using (3.5). To ease notation, define
$\hat{\beta}_{1,obs}$ $=$ $\text{logit}(\Pr(X_{2}=1|X_{1}=1,U=0))-\text{logit}(\Pr(X_{2}=1|X_{1}=0,U=0))$ (3.7)
$\hat{\beta}_{0,obs}$ $=$ $\text{logit}(\Pr(X_{2}=1|X_{1}=1,U=0))-\hat{\beta}_{1,obs}.$ (3.8)
Define the unobserved quantities $\hat{\beta}_{1,mis}$ and $\hat{\beta}_{0,mis}$ for the unit nonrespondents analogously, replacing $U_{i}=0$ with $U_{i}=1$ in (3.7) and (3.8). We set the log odds ratios equal for the unit respondents and nonrespondents, that is, we set $\hat{\beta}_{1,mis}=\hat{\beta}_{1,obs}.$ We also set
$\frac{\sum_{i\in\mathcal{S}}w_{i}I(x_{i1}^{*}=1,U_{i}=1)}{1+\exp(-(\hat{\beta}_{0,mis}+\hat{\beta}_{1,mis}))}+\frac{\sum_{i\in\mathcal{S}}w_{i}I(x_{i1}^{*}=0,U_{i}=1)}{1+\exp(-\hat{\beta}_{0,mis})}\\ =\hat{T}_{2}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i2}^{*}=1,U_{i}=0).$
To determine $\hat{\beta}_{0,mis}$, we numerically solve the system of nonlinear equations using the modified Powell method as implemented in the R package “nleqslv” (Hasselman, 2018). The imputation probabilities are then
$\Pr(X_{2}=1|x_{i1}^{*}=1,U_{i}=1)$ $=$ $\frac{1}{1+\exp[-(\hat{\beta}_{0,mis}+\hat{\beta}_{1,mis})]}$ (3.9)
$\Pr(X_{2}=1|x_{i1}^{*}=0,U_{i}=1)$ $=$ $\frac{1}{1+\exp(-\hat{\beta}_{0,mis})}.$ (3.10)
When $k>2$, we can apply the process used in (3.3) to each $X_{j}\in\mathcal{A}$. For example, if $X_{3}\in\mathcal{A}$, the analyst specifies a set of working probabilities, $p_{i3c}=\Pr(X_{i3}=c|x_{i1}^{*},x_{i2}^{*},w_{i})$ for $c=1,\dots,m_{3}$ via a (multinomial) logistic regression of $X_{3}$ on $(X_{1},X_{2})$ and possibly $W$. The analyst samples values of $(\hat{T}_{31},\dots,\hat{T}_{3m_{3}})$ and finds the values of $f_{3c}$ using expressions analogous to (3.3), replacing quantities based on $X_{2}$ with those based on $X_{3}$. When the working probabilities are identical across nonresponding units, analysts can solve a system of equations akin to those in (3.4) and (3.6).
@@ 3.3 Variables Without Margins for Unit Nonrespondents
After creating $L$ partially completed datasets using Section 3.1 and 3.2, we impute values $x_{ij}^{*}$ for unit nonrespondents for all $X_{j}\notin\mathcal{A}$. We use the random hot deck imputation procedure developed by Yang and Reiter (2025). We briefly summarize the procedure here, closely following their presentation.
For any unit $i\in\mathcal{S}$ with $U_{i}=0$, in any completed dataset, let $\boldsymbol{x}_{i}^{\mathcal{A}}=\{x_{ij}:X_{j}\in\mathcal{A};j=1,\dots,p;U_{i}=0\}$ be the values of the (possibly imputed) survey variables in the completed data for those $X_{j}\in\mathcal{A}$. Similarly, for any unit $i^{\prime}\in\mathcal{S}$ with $U_{i^{\prime}}=1$, let $\boldsymbol{x}_{i^{\prime}}^{\mathcal{A}*}=\{x_{i^{\prime}j}^{*}:X_{j}\in\mathcal{A};j=1,\dots,p;U_{i}=1\}.$ For each unit $i^{\prime}$ with $U_{i^{\prime}}=1$, in each completed dataset we construct its donor set, $\mathcal{D}_{i^{\prime}}=\{(x_{i1},\dots,x_{ip}):\boldsymbol{x}_{i}^{\mathcal{A}}=\boldsymbol{x}_{i^{\prime}}^{\mathcal{A}*},U_{i}=0,i\in\mathcal{S}\}$. We randomly sample one record $i$ from $\mathcal{D}_{i^{\prime}}$ and append its $\{(x_{i1},\dots,x_{ik}):X_{j}\notin\mathcal{A}\}$ to $\boldsymbol{x}_{i^{\prime}}^{\mathcal{A}*}$ to make the full imputation $\boldsymbol{x}_{i^{\prime}}^{*}$ for unit $i^{\prime}$. We apply this procedure for all units in $\mathcal{S}$ with $U_{i^{\prime}}=1$, resulting in a completed dataset for all $n$ units. We repeat this process in each of the $L$ datasets.
@@ 4 Simulation Studies
To demonstrate the methodology, we mimic the simulation study in Yang and Reiter (2025). They found that, by leveraging auxiliary margins, multiple imputation via the hybrid missingness MD-AM model can offer more accurate inferences than multiple imputation that does not utilize $\mathcal{A}$. However, they use (2.4) in lieu of the design weights for unit nonrespondents, whereas we use them per Section 3.
We generate $\mathcal{P}$ comprising $N=3,373,378$ units. This is the number of individuals in the 2022 American Community Survey data file used by Yang and Reiter (2025). We make one of the variables in this file a size variable $Z=(z_{1},\dots,z_{N})$; this is observed for all $N$ individuals. To generate a $\mathcal{S}$, we sample records from $\mathcal{P}$ independently with $\pi_{i}=1/10z_{i}$, where $i=1,\dots,N$. This Poisson sampling design results in approximately $n=6000$ units in any $\mathcal{S}$. Let $W=10Z$ be the survey weights.
For each unit $i=1,\dots,N$, we generate four binary variables $(x_{i1},x_{i2},x_{i3},x_{i4})$ and two continuous variables $(x_{i5},x_{i6})$. The data generation model is displayed in (4.1)–(4.5).
$U$ $\sim\text{Bernoulli}(\pi_{U})\quad\text{logit}(\pi_{U})=\nu_{0}$ (4.1)
$X_{1}|U$ $\sim\text{Bernoulli}(\pi_{x_{1}})\quad\text{logit}(\pi_{x_{1}})=\omega_{10}+\omega_{11}W+\theta_{1}U$ (4.2)
$X_{2}|X_{1},U$ $\sim\text{Bernoulli}(\pi_{x_{2}})\quad\text{logit}(\pi_{x_{2}})=\omega_{20}+\omega_{21}X_{1}+\theta_{2}U$ (4.3)
$X_{j}|X_{1},\dots,X_{j-1},U$ $\sim\text{Bernoulli}(\pi_{x_{j}})\quad\text{logit}(\pi_{x_{j}})=\omega_{j0}+\sum_{k=1}^{j-1}\omega_{jk}X_{k}\,\,\,\textrm{for }j=3,4$ (4.4)
$X_{j}|X_{1},\dots,X_{j-1},U$ $\sim\text{Normal}(\omega_{j0}+\sum_{k=1}^{j-1}\omega_{jk}X_{k},\sigma_{j}^{2})\,\,\,\textrm{for }j=5,6.$ (4.5)
We set $\nu_{0}=-1.2$ for a unit nonresponse rate around 23%. We consider $\theta_{1}\in\{-2,-0.5\}$. The marginal distribution of $X_{1}$ for unit respondents and nonrespondents is substantially different when $\theta_{1}=-2$ and modestly different when $\theta_{1}=-0.5$. We use $\theta_{2}=-2$, which ensures the unit respondents and nonrespondents have different marginal distributions of $X_{2}$. The remaining parameter values are available in Appendix A.
We independently sample 500 datasets. For each sample $\mathcal{S}$, we generate the unit nonrespondents by sampling Bernoulli random draws of $U_{i}$, where $i\in\mathcal{S}$, from the distribution of $U|X_{1},X_{2}$ implied by (4.1)–(4.3). We make $\boldsymbol{x}_{i}$ completely missing for any unit in $\mathcal{S}$ where the drawn $U_{i}=1$. For item nonresponse, we use the ICIN mechanisms
$R_{j}|X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},U=0\sim\text{Bernoulli}(\pi_{R_{j}}),\,\,\,\text{logit}(\pi_{R_{j}})=\phi_{j0}+\sum_{k\neq j}\phi_{jk}X_{k}$ (4.6)
@@
for $X_{j}$ where $j\in\{2,3,4,6\}$, and we let $X_{1}$ and $X_{5}$ be fully observed. We blank every $x_{ij}$ for which the sampled $R_{ij}=1$. We set parameters in (4.6) so that approximately 25% of each eligible survey variable is missing for units with $U_{i}=0$. Specific parameter values are available in Appendix A.
@@
We let $\mathcal{A}$ include $T_{1}=\sum_{i=1}^{N}x_{i1}$ and $T_{2}=\sum_{i=1}^{N}x_{i2}$, which we use for multiple imputation via the MD-AM models of Section 3. Using Section 3.1, we generate $L=10$ completed datasets for item nonresponse using a bespoke implementation of the “mice” package in R. Using Section 3.2, we consider two sets of working probabilities for $X_{1}$. The first allows $p_{i1c}$ to vary across units: each $p_{i1c}$ equals the unit’s predicted probability from the logistic regression of $X_{1}$ on $W$ estimated using the units with $U_{i}=0$. The second fixes $p_{i1c}=p_{1c}$ across units and uses (3.2). For both we use (3.9) and (3.10) to impute $X_{2}$. For the variances $V_{1}$ and $V_{2}$ used to sample $\hat{T}_{1}\sim N(T_{1},V_{1})$ and $\hat{T}_{2}\sim N(T_{2},V_{2})$, we use “mice” to make one completed dataset for all of $\mathcal{S}$ and compute the expressions for the unbiased variance estimators for $\hat{T}_{1}$ and $\hat{T}_{2}$ under Poisson sampling. Using Section 3.3, we create donor pools for the hot deck by matching on $(X_{1},X_{2})$.
@@
After generating $L=10$ completed datasets, we use the inferential methods in Rubin (1987) for the estimands in Yang and Reiter (2025), including marginal totals for $(X_{1},\dots,X_{6})$ and several probabilities involving multiple categorical variables. In each completed dataset, we compute Horvitz and Thompson (1952) point and variance estimators for a Poisson sampling design based on the design weights for all sampled records. For comparisons, we also use the imputation strategy in Yang and Reiter (2025) that does not utilize $\mathcal{A}$. It uses a bespoke implementation of “mice” to impute missing items and randomly samples whole records with replacement as imputations for unit nonresponse. We refer to this method as IH.
@@
Figure 4.1 summarizes the relative root mean squared errors (RMSEs) of the point estimates and empirical coverage rates of 95% multiple imputation confidence intervals for the three methods. Results in tabular form are at https://github.com/kevinxu47/MDAM. Both MD-AM models produce accurate point estimates, mitigating the bias from the nonignorable nonresponse evident in IH. The coverage rates for both MD-AM models tend to exceed 95% due to over-estimation of variances. As discussed by Yang and Reiter (2025), this is because the combining rules of Rubin (1987) do not account for using known totals in imputation. Overall, the results suggest that the MD-AM models enable analysts to incorporate known marginal totals in multiple imputation when using all sampled units’ design weights.
@@
Figure 4.1: Simulated relative RMSEs and coverage rates for population totals, two-way conditional probabilities, and joint probabilities for MD-AM and IH models. Left panel compares IH to the MD-AM model with fixed working probabilities. Right panel displays the two MD-AM models. Top row uses $\theta_{1}=-2$. Bottom row uses $\theta_{1}=-0.5.$
Appendix A Simulation Parameters
The parameter values used in the simulations of Section 4 are identical to those in Yang and Reiter (2025). We set $\nu_{u}=-1.2$; $(\omega_{10},\omega_{11})=(0.06,-0.0002)$ and $\theta_{1}\in(-2,-0.5)$; $(\omega_{20},\omega_{21})=(0.2,0.4)$ and $\theta_{2}=-2$; $(\omega_{30},\omega_{31},\omega_{32})=(0.2,0.3,0.1)$; $(\omega_{40},\dots,\omega_{43})=(0.2,0.4,0.4,0.1)$; $(\omega_{50},\dots,\omega_{54})=(0.4,1.2,-0.9,0.1,0.2)$ and $\sigma_{5}=0.5$; $(\omega_{60},\dots,\omega_{65})=(0.4,1.2,-0.9,0.1,-0.1,0.1)$ and $\sigma_{6}=0.5$; $(\phi_{j0},\dots,\phi_{j5})=(-1.4,0.1,0.1,0.1,0.1,0)$ for $j\in\{2,3,4\}$,; and, $(\phi_{60},\dots,\phi_{65})$$=(-1.4,0.1,0.1,0.1,0.1,0.1)$.

### v5

@@ 3.2.1 Multiplicative Adjustment Method
We impute each $x_{i2}^{*}$ given $x_{i1}^{*}$ and $\boldsymbol{z}_{i}$ (or $w_{i}$) for units with $U_{i}=1$ using a random draw from the adjusted probability mass function $\{\tilde{p}_{i2c}:c=1,\dots,m_{2}\}$, where $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}$ for $c=1,\dots,m_{2}-1$ and $\tilde{p}_{i2m_{2}}=1-\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}$. If $\sum_{c=1}^{m_{2}-1}f_{2c}{p}_{i2c}>1$, we compute (3.2) for $c=1,\dots,m_{2}$ and set $\tilde{p}_{i2c}=f_{2c}{p}_{i2c}/\sum_{c=1}^{m_{2}}f_{2c}{p}_{i2c}$. As for $X_{1}$, if some $f_{2c}{p}_{i2c}<0$, we set those $\tilde{p}_{i2c}=0$. This process ensures the completed-data Horvitz and Thompson (1952) estimates approximately match the sampled $(\hat{T}_{21},\dots,\hat{T}_{2m_{2}})$ in expectation, while also incorporating relationships between $X_{2}$, $X_{1}$ and $\boldsymbol{Z}$ (or $W$) implied in the working probabilities. We independently repeat the estimation and imputation process in each completed dataset. We refer to this method as MDAM–adj.
When $\{p_{i2c}\}$ derives from a logistic regression of $X_{2}$ on $X_{1}$ and $\boldsymbol{Z}$ (or $W$), using $\{\tilde{p}_{i2c}\}$ is equivalent to adjusting the intercept in that regression, leaving other coefficients alone. This implies a model in which the log-odds for $X_{1}$ are the same for unit respondents and nonrespondents, in accordance with no interactions between $U$ and $X_{1}$ in (2.2).
3.2.2 Systems of Equations Method
MDAM–adj captures the relationship between $X_{2}$ and $X_{1}$ through the logistic regression used in the working probabilities. Another approach is to impose constraints on the imputations through a system of equations implied by the assumptions underpinning (2.2) and (2.5), as we now describe.
As in Section 3.2.1, we first sample plausible values of $(\hat{T}_{21},\dots,\hat{T}_{2m_{2}})$. For now, we presume the working probabilities follow $p_{i2c}=p_{2cd}$ for all units. Within any completed dataset, we encode the constant log-odds assumption assumption in (2.2) as a set of $(m_{1}-1)(m_{2}-1)$ equations. For $c=2,\dots,m_{2}$ and $d=2,\dots,m_{1}$, we have
$\text{log}\left[\frac{p_{2cd}}{p_{21d}}\right]-\text{log}\left[\frac{p_{2c1}}{p_{211}}\right]=\text{log}\left[\frac{\Pr(X_{2}=c|X_{1}=d,U=0)}{\Pr(X_{2}=1|X_{1}=d,U=0)}\right]-\text{log}\left[\frac{\Pr(X_{2}=c|X_{1}=1,U=0)}{\Pr(X_{2}=1|X_{1}=1,U=0)}\right].$ (3.3)
We estimate the conditional probabilities for unit respondents in (3.3) via survey-weighted ratio estimators using the completed data; for example,
@@ 3.2.2 Systems of Equations Method
We encode the conditions on $(\hat{T}_{21}^{*},\dots,\hat{T}_{2m_{2}}^{*})$ in (2.5) as a set of $m_{2}$ equations. We have
$\sum_{d=1}^{m_{1}}\sum_{i\in\mathcal{S}}w_{i}{p}_{2cd}I(x_{i1}^{*}=d,x_{i2}^{*}=c,U_{i}=1)=\hat{T}_{2c}-\sum_{i\in\mathcal{S}}w_{i}I(x_{i2}^{*}=c,U_{i}=0).$ (3.5)
We solve these equations for the $m_{1}(m_{2}-1)$ conditional probabilities, $\{p_{2cd}:c=2,\dots,m_{2};d=1,\dots m_{1}\}$. We impute $x_{i2}^{*}$ given $x_{i1}^{*}$ for the unit nonrespondents using Bernoulli draws with probabilities $\{{p}_{2cd}\}.$ We refer to this method as MDAM–sys.
It is possible to modify MDAM–sys to use working probabilities $\{p_{i2c}\}$ that vary for units with the same value of $x_{i1}^{*}$. To do so, we impute using Bernoulli draws with probabilities $\{\tilde{p}_{i2c}\}$, where $\tilde{p}_{i2c}=f_{2cd}{p}_{i2c}$ and
$f_{2cd}=\frac{\sum_{i\in\mathcal{S}}w_{i}{p}_{2cd}}{\sum_{i\in\mathcal{S}}w_{i}{p}_{i2c}}$ (3.6)
for each $(X_{2}=c,X_{1}=d)$. The adjustment in (3.6) is motivated by matching $\sum_{i\in\mathcal{S}}w_{i}{p}_{2c}=\sum_{i\in\mathcal{S}}w_{i}f_{2cd}{p}_{i2c}.$
3.2.3 Accounting for Additional Variables with Margins
When $k>2$, for MDAM–adj we can apply the process used in (3.2) to each $X_{j}\in\mathcal{A}$. For example, if $X_{3}\in\mathcal{A}$, the analyst specifies a set of working probabilities, $p_{i3c}=\Pr(X_{i3}=c|x_{i1}^{*},x_{i2}^{*},\boldsymbol{z}_{i},w_{i})$ for $c=1,\dots,m_{3}$ via a logistic regression of $X_{3}$ on $X_{1},X_{2},$ and $\boldsymbol{Z}$ (or $W$). The analyst samples values of $(\hat{T}_{31},\dots,\hat{T}_{3m_{3}})$ and computes values of $f_{3c}$ using expressions analogous to (3.2), replacing quantities based on $X_{2}$ with those based on $X_{3}$. For MDAM–sys, the analyst can solve a system of equations akin to those in (3.3) and (3.5). However, MDAM–sys becomes increasingly complicated as more variables are added to $\mathcal{A}$.
@@ 3.3 Variables Without Margins for Unit Nonrespondents
After creating $L$ partially completed datasets using Section 3.1 and 3.2, we impute values $x_{ij}^{*}$ for unit nonrespondents for all $X_{j}\notin\mathcal{A}$. We use the random hot deck imputation procedure developed by Yang and Reiter (2025). Paraphrasing from their presentation, for any unit $i\in\mathcal{S}$ with $U_{i}=0$, in any completed dataset, let $\boldsymbol{x}_{i}^{\mathcal{A}}=\{x_{ij}:X_{j}\in\mathcal{A};j=1,\dots,p;U_{i}=0\}$ be the values of the (possibly imputed) survey variables in the completed data for those $X_{j}\in\mathcal{A}$. Similarly, for any unit $i^{\prime}\in\mathcal{S}$ with $U_{i^{\prime}}=1$, let $\boldsymbol{x}_{i^{\prime}}^{\mathcal{A}*}=\{x_{i^{\prime}j}^{*}:X_{j}\in\mathcal{A};j=1,\dots,p;U_{i}=1\}.$ For each unit $i^{\prime}$ with $U_{i^{\prime}}=1$, in each completed dataset we construct its donor set, $\mathcal{D}_{i^{\prime}}=\{(x_{i1},\dots,x_{ip}):\boldsymbol{x}_{i}^{\mathcal{A}}=\boldsymbol{x}_{i^{\prime}}^{\mathcal{A}*},U_{i}=0,i\in\mathcal{S}\}$. We randomly sample one record $i$ from $\mathcal{D}_{i^{\prime}}$ and append its $\{(x_{i1},\dots,x_{ik}):X_{j}\notin\mathcal{A}\}$ to $\boldsymbol{x}_{i^{\prime}}^{\mathcal{A}*}$ to make the full imputation $\boldsymbol{x}_{i^{\prime}}^{*}$ for unit $i^{\prime}$. We apply this procedure for all units in $\mathcal{S}$ with $U_{i^{\prime}}=1$, resulting in a completed dataset for all $n$ units. We repeat this process in each of the $L$ datasets.
@@ 4 Simulation Studies
We construct a population $\mathcal{P}$ comprising $N=3,405,809$ individuals from the 2023 American Community Survey public use files available from IPUMS. Each individual has the six variables described in Table 4.1. We fill in any missing values using a single bespoke run of the “mice” package. We treat the survey weight on the file as a known size measure, $Z=(z_{1},\dots,z_{N})$. To generate any $\mathcal{S}$, we sample records from $\mathcal{P}$ independently with probabilities $\pi_{i}=1/10z_{i}$, where $i=1,\dots,N$. This Poisson sampling results in approximately $n=6000$ individuals in any $\mathcal{S}$. We take 500 independent Poisson samples.
Table 4.1: Description of variables in the simulation study.
Variable Notation Range
Sex $X_{1}$ 1 = Male, 0 = Female
Marital status $X_{2}$ 1 = Married, 0 = Other
Any health insurance coverage $X_{3}$ 1 = Yes, 0 = No
School attendance $X_{4}$ 1 = Yes, 0 = No
Travel time to work $X_{5}$ 0 to 888
Occupational income score $X_{6}$ 0 to 80
For each $\mathcal{S}$, we generate unit nonresponse by sampling $U_{i}$ for all $i\in\mathcal{S}$ from
$U\sim\text{Bernoulli}(\pi_{U})\quad\text{logit}(\pi_{U})=\omega_{0}+\omega_{1}X_{1}+\omega_{2}X_{2}.$ (4.1)
We set $(\omega_{0},\omega_{1},\omega_{2})=(-1.6,0.5,0.5)$ for a unit nonresponse rate around 23%. We make all data other than $Z$ and $W$ completely missing for every unit where $U_{i}=1$. To generate item nonresponse for any unit $i$ with $U_{i}=0$, we sample $R_{ij}$ for $j=1,\dots,6$ using
$R_{j}\mid X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},Z,U=0\sim\mathrm{Bernoulli}(\pi_{R_{j}}),$ (4.2)
$\operatorname{logit}(\pi_{R_{j}})=\phi_{j0}+\sum_{k\neq j}\phi_{jk}X_{k}+\phi_{Z}Z.$
@@
We set parameters in (4.2) so that approximately 20%-25% of each survey variable is missing for units with $U_{i}=0$. Specifically, $\phi_{Z}=0.00001$; $\phi_{j0}=-1.6$ when $j\leq 4$ and $\phi_{j0}=-1.5$ when $j=5,6$; and, $\phi_{jk}=0.1$ when $k\leq 4$ and $\phi_{jk}=0.0001$ when $k=5,6$, for all $j$. We blank every $x_{ij}$ for which the sampled $R_{ij}=1$.
@@
We let $\mathcal{A}$ include $T_{X_{1}}=\sum_{i=1}^{N}x_{i1}$ and $T_{X_{2}}=\sum_{i=1}^{N}x_{i2}$. Using Section 3.1, we generate $L=10$ completed datasets for item nonresponse using a bespoke implementation of the “mice” package. Using Section 3.2, we implement MDAM–adj and MDAM–sys with working probabilities derived from logistic regressions of $X_{1}$ on $Z$ and of $X_{2}$ on $(X_{1},Z)$. For $V_{X_{1}}$ and $V_{X_{2}}$ used to sample $\hat{T}_{X_{1}}\sim N(T_{X_{1}},V_{X_{1}})$ and $\hat{T}_{X_{2}}\sim N(T_{X_{2}},V_{X_{2}})$, we use “mice” to make one completed dataset for all of $\mathcal{S}$ and compute the expressions for the unbiased variance estimators for $\hat{T}_{X_{1}}$ and $\hat{T}_{X_{2}}$ under Poisson sampling. Using Section 3.3, we create donor pools for the hot deck by matching on $(X_{1},X_{2})$.
@@
After generating $L=10$ completed datasets, we use the inferential methods in Rubin (1987) for the marginal totals for $(X_{1},\dots,X_{6})$ and several conditional and joint probabilities; see https://github.com/kevinxu47/MDAM for the full list. In each completed dataset, we compute Horvitz and Thompson (1952) point and variance estimators under Poisson sampling based on the design weights for all sampled records. For comparison, we also use the method in Yang and Reiter (2025) with the weights in (2.4). We refer to this method as MDAM–yr. Finally, we use a bespoke implementation of “mice” to impute missing items and randomly sample whole records with replacement as imputations for unit nonresponse. We refer to this method as IH. It does not utilize $\mathcal{A}$.
@@
Figure 4.1 summarizes the relative root mean squared errors (rRMSEs) of the point estimates and empirical coverage rates of 95% multiple imputation confidence intervals for the methods across the 500 samples. Results in tabular form are available at https://github.com/kevinxu47/MDAM. MDAM–adj tends to produce lower rRMSEs and higher coverage rates than IH. The coverage rates for the MD-AM models tend to exceed 95% due to over-estimation of variances. As discussed by Yang and Reiter (2025), this is because the combining rules of Rubin (1987) do not account for using known totals in imputation. MDAM–adj and MDAM–sys offer qualitatively similar results. For both, only the interval for $T_{X_{3}}$ has a lower than nominal coverage rate (around 85%). This rate approximately matches the coverage rate of design-based confidence intervals estimated before introducing any missing data (see the tabular results), suggesting that the lower than nominal coverage is a feature of the complete data estimator, not a consequence of MDAM modeling. MDAM–yr performs similarly to MDAM–adj and MDAM–sys. As evident in the tabular results, MDAM–yr tends to have lower variances, which results from flattening all weights for unit nonrespondents to a constant.
Overall, the results suggest that the new MD-AM models enable analysts to incorporate known marginal totals in multiple imputation when using all sampled units’ design weights.
@@
Figure 4.1: Simulated rRMSEs and coverage rates for population totals, two-way conditional probabilities, and joint probabilities. Top panel compares MDAM–adj and IH. Bottom panel displays MDAM–adj, MDAM–sys, and MDAM–yr.